OpenStudy (anonymous):

Factor completely: ax + 2ay + bx + 2by

OpenStudy (anonymous):

nevermind i got it now im so dumb *facepalm*

OpenStudy (anonymous):

ax+2ay+bx+2by =a(x+2y)+b(x+2y) =(a+b)+(x+2y)